OpenStudy (anonymous):

Solve, using the substitution method. m + n = 5 m – n = 3 A. There is no solution. B. The solution is (2, –1) C. The solution is (4, 1) D. There are an infinite number of solutions.

OpenStudy (anonymous):

C. The solution is (4, 1). Where: \[m = 4, n = 1\] Plug it in to see if it is true: \[4 + 1 = 5 \] \[4 – 1 = 3\] They are true, thus your answer must be C.

OpenStudy (anonymous):

Thanks! :D I have 4 more questions though plz :D

OpenStudy (anonymous):

These types pf questions are not that hard, attempt them if you don't understand them, post a new thread and tag me by doing: @LanguageEnthusiast I will try to help you if I can :-).

OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 A. (14, 12) B. (12, 14) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

@Wildberries What don't you understand? \[y = 4x – 6\] \[8x – 2y = 14\] You have to equations. Since the question is multiple choice simply plug in values that are given and try them out. Let us try the first answer choice: y = (4(14)) – 6 y = 50 Plug that into the second equation: \[8x – 2(50) = 14\] \[x = 14.250\] Thus A cannot be it, now try B yourself :-).

OpenStudy (sleepyjess):

@Wildberries ... do you understand how to do the substitution method? If not, I can explain it.

OpenStudy (anonymous):

yes